# Sheared flow profile effects on acoustic impedance eduction in small 3D ducts

### v1

@@ 4.1 Wavenumbers of the least-attenuated mode
As previously observed for the 2D duct (with a 1D flow profile) in Ref. [bonomoUnpublished], good agreement is found between the wavenumbers obtained under the uniform flow assumption and those from the exact solution of the PBE using the law-of-the-wall profile. This suggests that, at least for plane-wavelike modes, simplifying a 3D duct to a 2D or even uniform flow representation may not lead to significant inaccuracies. This contrasts with the findings of Ref. [roncen2020WavenumberBased], who considered a less physical hyperbolic tangent flow profile. The wavenumbers obtained considering the CFD flow profile agree well with those from both the IMBC and the tensorised law-of-the-wall profile, suggesting that the latter is a good representation for the duct geometry considered. Therefore, for the sake of conciseness, we adopt the tensorised law-of-the-wall profile as a representative realistic flow profile in the remainder of this work.
@@ 4.2 On the average Mach number definition
Figure 7: Least attenuated modes obtained for upstream propagation with the reference impedance at $3000\text{\,}\mathrm{Hz}$. (a) CHE+IMBC; (b) law-of-the-wall; (c) CFD, and; (d) hyperbolic tangent.
@@ 4.3 Simulated impedance eduction
In what follows, we apply the same procedure, but extend the analysis to include both the sliced and scaled velocity profiles for the 2D duct case, as well as the IMBC formulation using both $M$ and $M_{\mathrm{1D}}$ as input parameters. The numerical impedance eduction routine can be summarized as follows:
@@
    The PBE eigenvalue solver is used to compute the axial wavenumbers in the lined 3D duct with sheared grazing flow. A pair of axial wavenumbers is obtained, $k^{\pm}_{z,\mathrm{2D}}$, corresponding to the least attenuated modes. This step mimics a real-world impedance eduction approach where the wavenumbers are extracted from equally spaced acoustic pressure measurements using Prony-like algorithms [roncen2020WavenumberBased];
@@
We consider both the tensorised law-of-the-wall and hyperbolic tangent velocity profiles for data generation in this numerical experiment. The results are shown in Fig. 8.
@@
The analysis confirms key trends. First, consistent with Ref. [roncen2020WavenumberBased], the reduction to a 2D duct induces the classical mismatch between upstream and downstream propagation observed experimentally when the mean Mach number of the 1D profile, $M_{\mathrm{1D}}$, differs from the bulk Mach number of the 3D duct, $M$. This behavior holds for both velocity profiles considered. Second, when the 1D flow profile is scaled to match the original bulk Mach number, this discrepancy vanishes. This supports the findings of Ref. [jingInvestigationStraightforwardImpedance2015], who argued that averaging across the full cross-section yields the appropriate Mach number for eduction. Finally, impedances obtained via the traditional straightforward method, which assumes uniform flow and the Ingard–Myers boundary condition, remain a reasonable approximation, as long as the correct average Mach number is used, particularly when realistic flow profiles are employed to perform the in silico experiments.
@@ 5 Conclusions
This paper analyses the effects of two-dimensional (2D) sheared flow profiles on impedance eduction in three-dimensional (3D) ducts. The influence of adopting simplified velocity profile formulations or geometric simplifications—such as reducing the domain to 2D or assuming uniform flow with the Ingard–Myers Boundary Condition (IMBC)—was investigated. Wavenumbers obtained by solving the Pridmore–Brown equation (PBE) for different sheared flow profiles were compared to those from the Convected Helmholtz Equation (CHE) with IMBC. The results indicate that the IMBC offers a good approximation when compared to solutions based on realistic profiles, such as the universal law-of-the-wall or those derived from CFD simulations. Conversely, the widely used hyperbolic tangent profile yielded different wavenumbers. Additionally, it was shown that when reducing the domain to 2D, the average Mach number of the 1D profile must be scaled to match the bulk Mach number of the full 2D profile. A numerical experiment simulating traditional straightforward impedance eduction revealed no evidence that simplifying 3D duct acoustics to a 2D or uniform flow domain—when using realistic boundary layer representations—accounts for the directional discrepancies in educed impedance observed experimentally. This contrasts with the findings of Ref. [roncen2020WavenumberBased], who used a less realistic hyperbolic tangent profile. It is important to mention that viscous effects were neglected in this work and the investigation of the effect of viscosity in the same scenarios is suggested as future work.

### v3

@@ 3.2 Test Liner
Within this framework, the present work serves as a baseline study, providing a controlled reference against which more general formulations accounting for spatially varying and nonlinear impedance effects may be assessed in future work.
@@ 4.1 Wavenumbers of the least-attenuated mode
As a consistency check, wavenumbers generated for a prescribed mean-flow profile and impedance were used as input to the eduction procedure, which is described later in the manuscript. Under identical assumptions, the original impedance was recovered within numerical accuracy. It is observed that, for a given propagation direction, the axial wavenumbers exhibit limited sensitivity to the choice of mean flow profile. This suggests that the details of the shear profile alone are unlikely to explain the discrepancies observed between upstream and downstream impedance eduction, and reinforces the need to consider other factors, such as the effective mean flow Mach number used in the eduction model.
As previously observed for the 2D duct (with a 1D flow profile) in Ref. [bonomo2026Effect], good agreement is found between the wavenumbers obtained under the uniform flow assumption and those from the exact solution of the PBE using the law-of-the-wall profile. This suggests that, at least for plane-wavelike modes, simplifying a 3D duct to a 2D or even uniform flow representation may not lead to significant inaccuracies. This contrasts with the findings of Ref. [roncen2020WavenumberBased], who considered a less physical hyperbolic tangent flow profile. The wavenumbers obtained considering the CFD flow profile agree well with those from both the IMBC and the tensorised law-of-the-wall profile, suggesting that the latter is a good representation for the duct geometry considered. Therefore, for the sake of conciseness, we adopt the tensorised law-of-the-wall profile as a representative realistic flow profile in the remainder of this work.
@@ 4.2 On the average Mach number definition
Figure 7: Least attenuated modes obtained for upstream propagation with the reference impedance at $3000\text{\,}\mathrm{Hz}$. (a) CHE+IMBC; (b) law-of-the-wall; (c) CFD, and; (d) hyperbolic tangent. Color shading and contour lines represent the same normalized mode shape; contour lines are superimposed only to highlight spatial variations.
@@ 4.3 Simulated impedance eduction
In what follows, we apply the same procedure, but extend the analysis to include both the sliced and scaled velocity profiles for the 2D duct case, as well as the IMBC formulation using both $M$ and $M_{\mathrm{1D}}$ as input parameters. This numerical experiment mimics practical eduction setups in which the source location relative to the liner determines whether upstream- or downstream-propagating modes are identified and used in the eduction process. The numerical impedance eduction routine can be summarized as follows:
@@
    The PBE eigenvalue solver is used to compute the axial wavenumbers in the lined 3D duct with sheared grazing flow. A pair of axial wavenumbers is obtained, $k^{\pm}_{z,\mathrm{2D}}$, corresponding to the least attenuated modes. This step mimics a real-world impedance eduction approach where the wavenumbers are extracted from equally spaced acoustic pressure measurements using Prony-like algorithms [roncen2020WavenumberBased]. In summary, this is the only step that differs our in silico experiment from a real world experiment.
@@
It is important to emphasize that, in this numerical experiment, the axial wavenumbers computed from the 3D duct with a fully two-dimensional sheared mean flow are treated as reference data, representing an in silico impedance eduction experiment. The objective is to assess how the educed impedance is affected when this configuration is simplified to a two-dimensional duct model with a one-dimensional mean-flow profile. Accordingly, the impedance eduction in Eq. (18) is formulated as the minimization of the discrepancy between the reference axial wavenumbers obtained from the 3D configuration and those predicted by the reduced-order 2D model with a sliced or scaled velocity profile.
We consider both the tensorised law-of-the-wall and hyperbolic tangent velocity profiles for data generation in this numerical experiment. Fig. 8 summarizes the outcomes of the in silico eduction experiment for different flow representations and modeling assumptions. The figure compares the impedances educed from upstream- and downstream-propagating modes under varying levels of simplification, ranging from reduced-order 1D flow profiles to uniform-flow assumptions using the Ingard–Myers boundary condition. The purpose of this comparison is to assess whether the discrepancies between upstream and downstream eduction can be attributed to these modeling simplifications when the underlying reference data are generated from a realistic two-dimensional mean-flow profile.
@@
The analysis confirms key trends. First, consistent with Ref. [roncen2020WavenumberBased], the reduction to a 2D duct induces the classical mismatch between upstream and downstream propagation observed experimentally when the mean Mach number of the 1D profile, $M_{\mathrm{1D}}$, differs from the bulk Mach number of the 3D duct, $M$. This behavior holds for both velocity profiles considered. This observation is consistent with the NASA impedance eduction protocol [nark2018assessment], in which a bulk Mach number based on the duct mass flow rate is employed. The present results support the importance of using a consistent effective Mach number in the eduction model to ensure reliable impedance estimation. Second, when the 1D flow profile is scaled to match the original bulk Mach number, this discrepancy vanishes. This supports the findings of Ref. [jingInvestigationStraightforwardImpedance2015], who argued that averaging across the full cross-section yields the appropriate Mach number for eduction. Finally, impedances obtained via the traditional straightforward method, which assumes uniform flow and the Ingard–Myers boundary condition, remain a reasonable approximation, as long as the correct average Mach number is used, particularly when realistic flow profiles are employed to perform the in silico experiments. Importantly, these results indicate that the uniform-flow assumption and the Ingard–Myers boundary condition are not, by themselves, responsible for the upstream/downstream discrepancy, provided that the reference data are generated using a realistic mean-flow model.
These results further indicate that the observed upstream/downstream discrepancy is strongly linked to inconsistencies in the effective mean flow Mach number used in the eduction model, rather than to the specific choice of shear profile. When the effective Mach number is not preserved between the reference configuration and the reduced-order model, a mismatch between upstream and downstream propagation naturally arises. Conversely, enforcing this consistency leads to a collapse of the educed impedances, regardless of the level of flow profile simplification.
@@ 5 Conclusions
This paper analyses the effects of two-dimensional (2D) sheared flow profiles on impedance eduction in three-dimensional (3D) ducts, by generating data using an in silico acoustics model and then educing impedance using a simplified model. In particular, the influence of adopting simplified velocity profile formulations or geometric simplifications—such as reducing the domain to 2D or assuming uniform flow with the Ingard–Myers Boundary Condition (IMBC)—was investigated. It was found that simplifying the model used in the eduction process from 3D duct acoustics to a 2D or uniform flow domain does not account for the directional discrepancies in educed impedance observed experimentally, provided that data was generated from in silico experiments that used a realistic boundary layer representation. This contrasts with the findings of Ref. [roncen2020WavenumberBased], who used a less realistic hyperbolic tangent profile when generating their data. These findings show that the upstream/downstream discrepancies seen in experimental impedance eduction are not an inherent consequence of uniform-flow modeling or the Ingard–Myers boundary condition.
In addition, wavenumbers obtained by solving the Pridmore–Brown equation (PBE) for different sheared flow profiles were compared to those from the Convected Helmholtz Equation (CHE) with IMBC. The results indicate that the IMBC offers a good approximation when compared to solutions based on realistic profiles, such as the universal law-of-the-wall or those derived from CFD simulations, even though these profiles have significant sheared regions across much of the duct width. Conversely, the widely used hyperbolic tangent profile yielded different wavenumbers, even through the boundary layer thickness was matched to give the same boundary layer shape parameters. Additionally, it was shown that when reducing the domain to 2D, the average Mach number of the 1D profile must be scaled to match the bulk Mach number of the full 2D profile.
A key outcome of this study is that the consistency of the effective mean flow Mach number between the reference configuration and the eduction model is a primary factor in reducing the upstream/downstream discrepancy reported in the literature. This result suggests that previously observed differences attributed to shear flow effects may instead arise from inconsistencies in the eduction setup, particularly in the definition of the mean flow used in the propagation model. It is important to note that viscous effects were neglected in this work, and their influence under the same conditions remains to be investigated in future studies. Finally, the conclusions drawn here are based on two simplifying assumptions. First, a single duct geometry was considered; therefore, a parametric investigation of the effect of the duct cross-section is left for future work. Second, the mean flow was assumed to be fully developed and invariant along the duct axis. However, recent findings [paduano_impact_2025] suggest that the axial development of the boundary layer may introduce local effects on the liner impedance.